OpenStudy (anonymous):

find the volume generated by rotating the region enclosed by the graphs of y =4x^2 &y = 4x about the x-axis

OpenStudy (campbell_st):

start by finding where the graphs intersect...

OpenStudy (anonymous):

ok wait a sec i will

OpenStudy (anonymous):

(1,4)

OpenStudy (campbell_st):

well I would have thought (0,0) and (1, 4)  so to me it looks like the region is \[y = 4x - 4x^2\] so for volume you need to use \[V = \pi \times \int\limits_{0}^{1} y^2 dx\] so you need to go and find the value of \[y^2 ...if... y = 4x - 4x^2\] then substitute and evaluate the definite integral. hope it helps...